### v2

# How Similar Are Grokipedia and Wikipedia? A Multi-Dimensional Textual and Structural Comparison

Abstract
Abstract
The launch of Grokipedia, an AI-generated encyclopedia developed by Elon Musk’s xAI, was presented as a response to perceived ideological and structural biases in Wikipedia, aiming to produce “truthful” entries via the large language model Grok. Yet whether an AI-driven alternative can escape the biases and limitations of human-edited platforms remains unclear. This study undertakes a large-scale computational comparison of 1,800 matched article pairs between Grokipedia and Wikipedia, drawn from the 2,000 most-edited Wikipedia pages. Using metrics across lexical richness, readability, structural organization, reference density, and semantic similarity, we assess how closely the two platforms align in form and substance. The results show that while Grokipedia exhibits strong semantic and stylistic alignment with Wikipedia, it typically produces longer but less lexically diverse articles, with fewer references per word and greater structural variability. These findings suggest that AI-generated encyclopedic content currently mirrors Wikipedia’s informational scope but diverges in editorial norms, favoring narrative expansion over citation-based verification. The implications highlight new tensions around transparency, provenance, and the governance of knowledge in an era of automated text generation.
1 Introduction
Online encyclopedias have become infrastructural to public knowledge, shaping how people learn about politics, science, culture, and current events. Wikipedia, launched in 2001, remains the paradigmatic example: a nonprofit, volunteer-edited project governed by policies such as Neutral Point of View (NPOV) and verifiability. Despite its success, Wikipedia has long faced questions about bias, reliability, and systemic underrepresentation. [^1] For a survey of evidence and debates on reliability and bias, see e.g., [4, 8, 17, 18].
In late 2025, xAI introduced Grokipedia, an AI-generated encyclopedia positioned as an explicit alternative to Wikipedia. [^2] See [20, 1, 19, 11, 13, 6, 10]. According to xAI’s framing, Grokipedia aims to “purge out the propaganda” and provide “truthful” entries, with content generated and internally “fact-checked” by the Grok language model rather than curated by a community of human editors.[20, 11, 10] At launch (October 27, 2025), Grokipedia reportedly contained roughly 800–900 thousand entries, supported a suggest-edit rather than direct-edit workflow, and offered only limited transparency about licensing and code provenance.[20, 19, 1] Early commentary focused on two recurring issues: (i) claims that many Grokipedia pages were copied or closely adapted from Wikipedia, and (ii) concerns about ideological bias, hallucinated citations, and other generative artifacts typical of large language models.[13, 6, 11, 10]
The launch has reignited a longstanding question: is Wikipedia biased, and if so, can an AI-first encyclopedia do better? Empirically, prior research documents multiple bias forms in Wikipedia, early left–right slant in political pages that attenuated over time,[4] topical and coverage imbalances,[8] and framing asymmetries across ideology and gender.[17, 18] The dynamics of editorial conflict and consensus formation have also been studied extensively, showing that edit activity, controversy, and attention are strongly correlated.[26] Parallel work on large language models (LLMs) shows measurable political and cultural biases that vary across model families and prompts,[22, 15, 3, 9] raising distinct questions about the epistemic reliability of AI-generated text.
Against this backdrop, our study asks a concrete, testable question: for topics that exist on both platforms, how similar are Grokipedia and Wikipedia in practice? We address this by pairing articles across platforms and comparing them along four analytic dimensions: (1) lexical similarity (TF–IDF cosine, unigram and $n$-gram overlap), (2) semantic similarity (embedding cosine and BERTScore), (3) structural similarity (section and paragraph organization), and (4) stylistic similarity (readability, lexical diversity, part-of-speech composition). In addition, we examine descriptive indicators, article length, readability, reference density, and link structure, since editorial workflows (human versus model-generated) shape article form as much as content.
To construct a broad and representative corpus, we began with Wikipedia’s 2,000 most-edited English-language articles, a population empirically linked to controversy and social salience.[26] Among these, approximately 1,800 titles had valid, nontrivial matches on Grokipedia. This enlarged sample enables more stable estimates of cross-platform differences than earlier pilot comparisons limited to a few hundred pairs.
This comparison provides empirical grounding for evaluating the distinctiveness, or sameness, of Grokipedia relative to Wikipedia. If Grokipedia largely reproduces Wikipedia’s textual and structural patterns, the platform may represent an AI-mediated continuity rather than a substantive editorial departure. Conversely, systematic differences in length, style, or citation density could reveal an emerging form of automated editorial logic, distinct from Wikipedia’s norms of transparency, consensus, and verifiability. By quantitatively characterizing these dimensions across matched topics, we aim to clarify where the two encyclopedias align, where they diverge, and what those patterns imply for neutrality, provenance, and the evolving governance of reference knowledge in the age of generative AI. [^3] For broader debates on neutrality, transparency, and provenance in AI-generated reference content, see [11, 6, 10, 23].
2 Methods
2.1 Sampling and Data Collection
We analyzed the 2,000 most-edited English-language Wikipedia articles as of October 2025, identified via cumulative edit counts. Prior research shows that heavily edited entries correlate strongly with controversy, topical salience, and social polarization [26]. This sampling strategy thus prioritizes articles that are both textually substantial and socially significant.
For each title, we retrieved the corresponding entries from Wikipedia and Grokipedia, generating URLs of the form https://en.wikipedia.org/wiki/<Title>
and
https://Grokipedia.com/page/<Title>. HTML pages were downloaded via the requests library (Python 3.11) with polite delays and a standard user-agent header, then parsed using BeautifulSoup4. The extraction and analysis pipeline continuously wrote progress files (wiki_Grokipedia_similarity_live.csv) to enable incremental recovery and monitoring. After host-aware text extraction (see below), only pairs in which both pages contained at least 500 words of clean prose were retained. Of the 2,000 target titles, 1,811 matched article pairs satisfied these criteria and formed the final analytical sample.
2.2 Text Cleaning and Feature Extraction
We implemented a host-aware parsing strategy tailored to each platform’s HTML structure to maximize content fidelity.
For Wikipedia, extraction was restricted to the #mw-content-text container and limited strictly to <p> and <li> elements. Elements such as infoboxes, hatnotes, metadata, tables, navboxes, and reference lists were removed prior to text collection.
For Grokipedia, a more adaptive extractor identified the main article-like container (<main>, <article>, or the largest <div>/<section>) and retained its narrative text while filtering out menus, advertisements, and high link-density regions. Scripts, styles, and sidebars were removed in both cases.
Each cleaned article was tokenized into sentences and words using nltk’s Punkt tokenizer. A comprehensive set of descriptive, structural, and stylistic metrics was computed:
  • Structural features: counts of paragraphs, headings (h1–h4), hyperlinks, images, and references; visible character counts; and derived ratios such as references, links, and headings per 1,000 words.
  • Stylistic and readability features: average sentence length, lexical diversity (type–token ratio), Flesch–Kincaid grade level, Gunning–Fog index, and part-of-speech composition (nouns, verbs, adjectives, adverbs) using spaCy’s en_core_web_sm model.
  • Lexical density and reading time: ratio of unique to total words and estimated reading time assuming 200 words per minute.
All features were computed independently for both platforms and labeled with prefixes a_ (Grokipedia) and b_ (Wikipedia).
2.3 Similarity Measures Between Platforms
To quantify cross-platform alignment, we computed a suite of nine similarity measures grouped into four conceptual domains:
  1. 1.
    Lexical similarity: cosine similarity of TF–IDF vectors (1–2 grams, English stop-words removed) and unigram Jaccard index.
  2. 2.
    N-gram overlap: overlap coefficients for 1-, 2-, and 3-gram sequences, capturing local phrase reuse.
  3. 3.
    Semantic similarity: cosine similarity between SentenceTransformer embeddings (all-MiniLM-L6-v2[14], and contextual similarity via BERTScore F1 for the first 50 sentences per article [27].
  4. 4.
    Structural and stylistic similarity: a composite of (a) overlap in detected section headings and (b) proportional similarity of median paragraph lengths (weighted 0.6 and 0.4, respectively), together with normalized Manhattan distance across stylistic profiles (sentence length, lexical diversity, readability, POS composition) transformed to a 0–1 scale.
A composite similarity index was calculated as the mean of lexical, semantic, structural, and stylistic sub-scores, providing a general indicator of cross-platform convergence.
2.4 Statistical and Comparative Analysis
All analyses were performed in Python 3.11 using pandas, numpy, and scikit-learn. Descriptive statistics (mean $\pm$ SD) were calculated for all features. Differences between Grokipedia and Wikipedia were evaluated using paired $t$-tests for normally distributed measures and Wilcoxon signed-rank tests otherwise. Spearman rank correlations among similarity metrics were computed to assess interdependence and clustering. Visualizations, including histograms, correlation heatmaps, and mean $\pm$ SE summary plots, were generated with matplotlib and exported directly to PDF for inclusion in the manuscript.
2.5 Limitations
The dataset covers only topics that are both heavily edited on Wikipedia and available on Grokipedia. Because edit frequency correlates with controversy and visibility [26], the sample likely overrepresents politically or culturally salient topics. In addition, Grokipedia’s content-generation pipeline is opaque, potentially introducing stylistic inflation (e.g., verbosity, hallucinated references) or structural artifacts that differ from human-edited prose. Nevertheless, the scale of the dataset and the matched-pair design provide a robust basis for assessing how human- and AI-generated encyclopedic text diverge in length, structure, and informational alignment.
3 Results
3.1 Sample and Coverage
From the expanded list of the 2,000 most-edited English Wikipedia titles, we successfully retrieved and matched 1,811 Grokipedia–Wikipedia article pairs that passed quality checks (both pages available and $\geq$500 cleaned words each). [^4] Clean word counts were computed from visible prose only, excluding infoboxes, scripts, and navigation templates. This larger corpus substantially extends the coverage of the original 416-title sample while preserving topical breadth and diversity.
3.2 Descriptive Differences Between Platforms
Figure 1 compares article-length distributions on a logarithmic scale. Grokipedia entries remain systematically longer than their Wikipedia counterparts, though the gap narrows compared with the smaller sample. Readability, measured by the Flesch–Kincaid grade level (panel B of Figure 2), is again higher on Grokipedia, indicating more syntactically complex and less reader-friendly prose.
Table 1 summarises key descriptive statistics. Across the enlarged dataset, Grokipedia articles average about 14,200 words versus 9,400 on Wikipedia, and they exhibit lower lexical diversity and far fewer explicit references, links, and headings per thousand words. These contrasts reinforce that Grokipedia expands text length through elaboration rather than citation density or structural detail. Figure 2 visualises these platform differences (means $\pm$ SE) across all six measures reported in Table 1.
Figure 1: Distribution of clean word counts by platform (log scale).
Figure 2: Platform differences for key descriptives (means $\pm$ standard error). Panels show: (A) clean word count (log10), (B) Flesch–Kincaid grade, (C) lexical diversity (TTR), (D) references per 1k words, (E) links per 1k words, and (F) headings (H2–H4) per 1k words.
Table 1: Descriptive statistics across matched article pairs (N=1,811N{=}1{,}811). Values are means [SD]; $p$ from paired $t$-tests.
Measure Grokipedia Wikipedia $\Delta$ $p$
Clean word count 14,200 [5,650] 9,360 [4,710] +4,840 $10^{-200}$
Flesch–Kincaid grade 19.3 [6.6] 12.8 [9.9] +6.5 $10^{-100}$
Lexical diversity (TTR) 0.244 [0.057] 0.248 [0.049] -0.004 0.02
References per 1k words 21.4 [7.4] 90.2 [218] -68.8 $10^{-50}$
Links per 1k words 16.8 [6.0] 428 [683] -411 $10^{-120}$
Headings (H2–H4)/1k words 2.87 [1.20] 5.58 [8.32] -2.71 $10^{-40}$
3.3 Cross-Platform Similarity
To assess how closely Grokipedia mirrors Wikipedia, we computed lexical, semantic, structural, and stylistic similarity metrics for each article pair. Figure 3 shows the correlation structure among these metrics: lexical and semantic measures form a tight cluster, while structural and stylistic dimensions vary more independently. Table 2 reports the average values across all pairs.
Figure 3: Correlation among similarity metrics (Pearson’s $r$) across 1,811 article pairs.
Table 2: Average similarity metrics across matched pairs (N=1,811N{=}1{,}811).
Metric Mean [SD]
TF–IDF cosine 0.505 [0.149]
Jaccard (unigram) 0.362 [0.217]
Overlap (1-gram) 0.597 [0.213]
Overlap (2-gram) 0.329 [0.339]
Overlap (3-gram) 0.250 [0.374]
Semantic cosine 0.825 [0.106]
BERTScore (F1) 0.193 [0.334]
Structural similarity 0.216 [0.163]
Stylistic similarity 0.835 [0.072]
3.4 Distribution of Similarity Scores
Figure 4 compiles the histograms for all nine similarity metrics. Most distributions are unimodal and right-skewed, especially lexical and semantic measures, indicating that the majority of Grokipedia pages are moderately to highly aligned with their Wikipedia equivalents. Structural similarity, by contrast, is more dispersed, reflecting Grokipedia’s simpler sectioning and reference patterns.
Figure 4: Empirical distributions of similarity scores across metrics. Each panel corresponds to one of the nine metrics.
Together, these findings reinforce that Grokipedia functions less as an independent encyclopedic corpus than as an AI-generated echo of Wikipedia, maintaining meaning and tone while attenuating the structural and evidentiary scaffolding that underpins human-curated knowledge.
4 Discussion and Conclusion
This study presents the first large-scale, systematic comparison between Grokipedia and Wikipedia across roughly 1,800 matched article pairs drawn from the 2,000 most-edited English Wikipedia titles. Despite their distinct production paradigms, community-driven versus AI-generated, the two encyclopedias exhibit remarkable alignment at the level of meaning and style. Semantic similarity averages around 0.82 and stylistic similarity around 0.83, indicating that Grokipedia reproduces much of Wikipedia’s linguistic and conceptual structure. However, the platforms diverge substantially in form and informational scaffolding. Grokipedia articles are, on average, several times longer, with higher Flesch–Kincaid grade levels (greater syntactic complexity) but lower lexical diversity and reference density. This pattern suggests that Grokipedia’s generation process elaborates on existing material, expanding text length and rhetorical flow, rather than producing substantively new or more rigorously sourced knowledge.
The similarity distributions reveal further nuance. While semantic and lexical measures are unimodal, reflecting a broad alignment across most topics, structural and stylistic similarities are more dispersed, suggesting heterogeneous modes of generation. Some Grokipedia pages closely mirror Wikipedia’s organization and tone; others diverge sharply, producing verbose or restructured narratives. This heterogeneity implies a hybrid authorship logic: part replication, part improvisation.
These results also resonate with earlier linguistic analyses of Wikipedia itself. Yasseri, Kornai, and Kertész [2012a] showed that language complexity in Wikipedia varies systematically with editorial dynamics: simplification efforts reduce syntactic depth without necessarily affecting lexical richness, and conflict intensity correlates with readability shifts. Our findings parallel this pattern in reverse, Grokipedia’s language inflates sentence length and readability grade without a proportional increase in lexical variety. In both cases, surface-level stylistic change (simplification or elaboration) alters perceived accessibility more than underlying vocabulary structure. Moreover, since our corpus also centers on heavily edited and hence often controversial topics, the observed stylistic divergence may partly reflect how both human and AI systems respond to socially contested knowledge: by expanding exposition rather than diversifying substance.
Elon Musk framed Grokipedia as an antidote to Wikipedia’s “propaganda” and ideological bias [2, 7, 24]. Yet early independent assessments from outlets including The Verge, Wired, The Washington Post, and TIME found that many Grokipedia articles are derived from Wikipedia, often copied or paraphrased, while selectively emphasizing Musk’s personal and political narratives [16, 12, 5, 21]. Our quantitative results support this view: Grokipedia achieves high semantic alignment with Wikipedia but introduces stylistic inflation and reduced citation density. In other words, Grokipedia’s output largely repackages existing human-curated content through an AI lens that privileges fluency and narrative over attribution. Rather than “purging bias,” the system appears to re-encode it, subtly and opaquely, within model parameters and prompt conditioning.
As argued by Yasseri [24], Wikipedia’s openness transforms bias from a flaw into a form of epistemic visibility: disputes, edits, and talk pages make disagreement transparent and correctable. Grokipedia inverts this model. Its authorship is singular, automated, and invisible, bias becomes latent, hidden within model weights and unseen editorial heuristics. This creates what might be termed an epistemic opacity paradox: Grokipedia appears neutral because it lacks human editors, yet its underlying generative logic is uninspectable and unaccountable. The contrast highlights a deeper epistemological divide between collective knowledge systems, where disagreement is traceable, and algorithmic knowledge systems, where authority is inferred but not negotiated.
Several limitations qualify these findings. First, the dataset focuses on Wikipedia’s 2,000 most-edited pages, of which around 1,800 have counterparts on Grokipedia. This selection likely overrepresents high-profile, contentious topics, those most prone to editing wars or ideological scrutiny [26]. Second, the similarity metrics employed, lexical, semantic, structural, and stylistic, quantify form and textual alignment but not factual accuracy or ideological framing. The presence of hallucinated claims, selective omissions, or subtle rhetorical shifts remains outside the scope of automated comparison. Third, both platforms are dynamic: Wikipedia continuously evolves, while Grokipedia’s generative parameters may change with future model retraining. Finally, Grokipedia’s underlying data sources and editorial interventions are opaque, preventing full provenance auditing or causal inference about model bias.
Despite its stated aim to “correct” Wikipedia’s ideological slant, Grokipedia currently functions less as an epistemic alternative and more as a synthetic derivative of Wikipedia. It mirrors Wikipedia’s informational scope and linguistic tone but replaces community deliberation with algorithmic synthesis. The result is an encyclopedia that is fluent yet fragile, expansive in form but shallow in verifiability. In effect, Grokipedia trades collective accountability for computational authority.
These findings extend beyond a single platform. As large language models increasingly mediate the production of knowledge, they risk conflating eloquence with accuracy and replacing transparency with automation. The challenge ahead is therefore not only technical, improving factuality or retrieval, but institutional: ensuring that provenance, bias, and contestation remain visible in an age of synthetic knowledge. Future research should move beyond surface similarity to assess factual divergence, ideological asymmetry, and user trust. Comparing how readers evaluate credibility across human-edited and AI-generated sources may reveal whether Grokipedia’s expansion of text corresponds to an expansion, or erosion, of understanding.
5 Acknowledgement
This publication has emanated from research supported in part by grants from Taighde Eireann-Research Ireland under Grant numbers IRCLA/2022/3217. We acknowledge support from Workday, Inc.